OpenStudy (theanarchist):

If your right then you get a medal. 60+60-40+75+39X0+1=?

OpenStudy (theanarchist):

Your wrong too my friend

OpenStudy (anonymous):

the way she wrote the equation is wrong it should be 60+60-40+75+39(0+1)= then it would be 194

OpenStudy (theanarchist):

Nope sorry

OpenStudy (anonymous):

oh i got it it's 81

OpenStudy (mathstudent55):

156

OpenStudy (theanarchist):

Its 1 everybody...What ever all those numbers equal too then its multiplied by zero that number now equals zero (ANYTHING multiplied by 0 is 0) so its 0 plus 1then that is 1. So like I said its 1 everybody *mind blown*

OpenStudy (mathstudent55):

60 + 60 - 40 + 75 + 39 X 0 + 1 = 156 (60 + 60 - 40 + 75 + 39) X 0 + 1 = 1 Parentheses make a big difference.

OpenStudy (theanarchist):

No not really that really just a mind thing in this situation. Type it into a calculator with out parentheses to make sure of yourself. Nice try though @mathstudent55

OpenStudy (anonymous):

it do pemdas parentheses first

OpenStudy (anonymous):

@mathstudent55 without the parentheses it is 156 but with it it's 81

OpenStudy (theanarchist):

If I did type it into parentheses then it would be real easy for you to copy and paste it. btw

OpenStudy (mrnood):

@TheAnarchist the way you wrote it, in normal maths notation does NOT evaluate to 1 you need to follow the pedmas rules or use parentheses Without them it is 156

OpenStudy (anonymous):

right @MrNood

OpenStudy (mrnood):

and whilst in 'pedantry' mode it's "you're wrong" not "your wrong"

OpenStudy (anonymous):

@TheAnarchist it was'nt a copypaste thing i don't do that BTW i was checkin my answer

OpenStudy (theanarchist):

Hey smart one were in the Mathematics section not the English section D-Bag

OpenStudy (anonymous):

dang

OpenStudy (mrnood):

Precise communication is a general requirement for educated people - it does not follow 'subjects' The real world is not made up of subjects

OpenStudy (mrnood):

maybe "your wrong" is an equivalent of "my bad" That would work

OpenStudy (mathstudent55):

The way I placed the parentheses, it does evaluate to 1. Since the original expression in the post evaluates to 156, and I have no way of knowing if parentheses were meant to be there, and if so, where the parentheses should go, I placed parentheses where I felt like, and it does evaluate to 1 the way I placed the parentheses. If it is supposed to evaluate to 81, then show me where the parentheses go.

OpenStudy (mrnood):

and BTW "we're" in the Maths section - maybe you "were" in the English section

OpenStudy (mathstudent55):

The fact that if you enter this calculation in the order written into a calculator, you don't get the correct result does not mean anything because most simple calculators are not algebraic calculators and do not follow the correct order of operations, known to many English-speaking students as PEMDAS. Incidentally, most people who know of PEMDAS don't really know how to use it and explain it incorrectly.